OpenStudy (anonymous):

Will fan and medal. A design on the surface of a balloon is 5 cm wide when the balloon holds 71 cm of air. How much air does the balloon hold when the design is 10 cm wide? Explain the method you use to find the amount of air.

OpenStudy (anonymous):

I know the answer i'm just not sure how to explain the method

OpenStudy (md152727):

This question is related to how volume changes when you change the size of something. Think of a cube that is 1cm by 1 cm by 1cm. It will have volume of 1*1*1=1cm^3, but if you double the size of the cube, the volume will now be 2*2*2=8 and if you triple it it wll be 3*3*3=27. So changing all the linear dimensions of something by a factor will change the volume by that factor cubed. This works for all 3d shapes as long as all dimensions are changing at the same rate. for a balloon, this is true. if you double the width of the balloon, its volume will change by a factor of 2*2*2=8. The inital volume was 71cm so the final will be 8 times 71 i,e, 568 cm^3

OpenStudy (anonymous):

i think 76

OpenStudy (anonymous):

So what would the method be called? I have that answer im just not sure how to explain it

OpenStudy (anonymous):

No, it's 568